Solve the quadratic equation by completing the square.
2x^2+10x+11=0
Thanks in advance.

Answer by Nate(3500)

2x^2 + 10x + 11 = 0
2x^2 + 10x = -11
x^2 + 5x = -11/2
(x + 5/2)^2 = -11/2 + 25/4 = -22/4 + 25/4 = 3/4
x + 5/2 = +- sqrt(3)/2
x = (-5 +- sqrt(3))/2